PHASE SHIFT CAVITY RING DOWN (PS-CRD) SPECTROSCOPY AND OVERTONE SPECTRA.

\underline{CARLOS MANZANARES I.}, ERNEST K. LEWIS, Department of Chemistry and Biochemistry, Baylor University, Waco, TX 76798; and MARY C. SALAZAR, ANTONIO J. HERNANDEZ, Department of Chemistry, Simon Bolivar University, Caracas 1080A, Venezuela.

Phase Shift Cavity Ring Down (PS-CRD) spectroscopy was used for the detection of (v = 0 → 6) C-H vibrational transition of deuterated methanes and the (v = 0 → 5) transition of HD at room temperature. A continuous wave dye laser in the wavelength range from 600 to 650nm and with a 0.5 cm^{−1} resolution was electro-optically modulated, and passed through an optical cavity. The change in the phase angle of the modulated laser beam after passing through the cavity was recorded as a function of the wavelength. Absolute cross sections were obtained.